Consider the motion of a particle along a spiral given by r=R-Aθ [read as R minus A(theta)].The position of particle at time t=0 is r=R and θ=0. The speed of particle decreases linearly with time as v(t) =u - kt from time t=0 to t=u/k. What is the velocity as function of time in polar coordinates?


1
Expert's answer
2021-10-01T12:21:27-0400

We have to use the definitions for r and v to find the velocity as a function of time in polar coordinates


"r=R-A\theta; R, A=cte
\\ v= u-kt; u,k=cte
\\ \text{we also have } v=\cfrac{dr}{dt}=\cfrac{d}{dt}(R-A\theta)=-A\cfrac{d\theta}{dt}
\\ \text{which implies that } -A\cfrac{d\theta}{dt}=u-kt
\\ \implies \text{and finally } \cfrac{d\theta}{dt}=\cfrac{kt-u}{A}"


In conclusion, we find that "\cfrac{d\theta}{dt}=\cfrac{kt-u}{A}", where k, u and A are constants.


Reference:

  • Sears, F. W., & Zemansky, M. W. (1973). University physics.